Grade 11Organic Chemistry

What will be normality of resulting solution when 100 ml of 0.5 normal solution of hcl is mixed with 200 ml of 0.5 normal of hcl

Mehrab
8 Years agoGrade 11

1 Answer

Mahima Kanawat
8 Years ago
Dear student... N1V1+N2V2=NV ; 100×0.5+200×0.5=N×300 ; N=0.5... WITH REGARDS, MAHIMA (ASKIITIANS FORUM EXPERT)